# Diffusion Beats Autoregressive in Data-Constrained Settings

### v3

@@ 1 Introduction
LLM development has so far been dominated by autoregressive (AR) models, which factorize the joint distribution of text in a fixed left-to-right order, i.e., $p(x)=p(x_{0})p(x_{1}\mid x_{0})p(x_{2}\mid x_{1})p(x_{3}\mid x_{1},x_{2})$. While this modeling approach has delivered state-of-the-art performance across a range of benchmarks, it remains unclear whether it is the optimal strategy going forward. Recently, diffusion-based models—specifically masked diffusion models  austin2021structured ; sahoo2024simple ; lou2310discrete ; shi2024simplified ; arriola2025block —have emerged as an alternative strategy, where they model text via a random-order factorization, e.g., $p(x)=p(x_{2}|x_{3})p(x_{3}|x_{0})p(x_{0}|x_{1})p(x_{1})$. These models learn across many different masked versions of the text, allowing them to train on diverse token orderings. Although diffusion language models have demonstrated similar scaling behavior to AR models (nie2024scaling, ; swerdlow2025unidisc, ), their practical benefits have, so far, been modest—largely due to their high training compute requirements.
@@ 4.1 Does Diffusion Beat AR in Data-Constrained Settings?
Prior comparisons between diffusion and autoregressive (AR) language models have largely focused on the single-epoch regime, where each token is seen only once during training mdlm ; nie2024scaling ; swerdlow2025unidisc . In this setting, diffusion models are consistently reported to require substantially more training compute ($C\sim 6ND$) than AR models to achieve comparable validation loss. For instance, Nie et al. (nie2024scaling, ) and Swerdlow et al. (swerdlow2025unidisc, ) derive scaling laws showing that masked diffusion models can require up to 16$\times$ more compute than AR counterparts.

### v4

@@ 1 Introduction
LLM development has so far been dominated by autoregressive (AR) models, which factorize the joint distribution of text in a fixed left-to-right order. While this modeling approach has delivered state-of-the-art performance across a range of benchmarks, it remains unclear whether it is the optimal strategy going forward. Recently, diffusion-based models—specifically masked diffusion models  austin2021structured ; sahoo2024simple ; lou2310discrete ; shi2024simplified ; arriola2025block —have emerged as an alternative strategy, framing text generation as an iterative masked denoising process rather than next-token prediction. At each step, the model predicts a randomly masked subset of tokens conditioned on the remaining ones, implicitly averaging over many conditional prediction orders instead of committing to one. Although these models have demonstrated similar scaling behavior to AR models (nie2024scaling, ; swerdlow2025unidisc, ), their practical benefits have, so far, been modest—largely due to their high training compute requirements.
@@ 4.1 Does Diffusion Beat AR in Data-Constrained Settings?
Prior comparisons between diffusion and autoregressive (AR) language models have largely focused on the single-epoch regime, where each token is seen only once during training nie2024scaling ; swerdlow2025unidisc . In this setting, diffusion models are consistently reported to require substantially more training compute ($C\sim 6ND$) than AR models to achieve comparable validation loss. For instance, Nie et al. (nie2024scaling, ) and Swerdlow et al. (swerdlow2025unidisc, ) derive scaling laws showing that masked diffusion models can require up to 16$\times$ more compute than AR counterparts.